E until K please!
Thanks!

Saxbury Corporations relevant range of activity is 3,000 units to 7,000 units. When it produces and sells 4,000 units, its average costs per unit are as follows: 6.80 $3.50 1.60 $4.00 $0.75 s 0.60 s 0.50 $0.50 Direct labor Fixed nanufacturing overhead Fixed selling expense Fixed ve expense a. For financial reporting purposes, what is the total amount of product costs incurred to make 4,000 units? b. For financial reporting purposes, what is the total amount of period costs incurred to sell 4,000 units? c. If 5,000 units are sold, what is the variable cost per unit sold? (Round Per unit answer to 2 decimal places.) d. If 5,000 units are sold what is the total amount of variable costs related to the units sold e. If 5,000 units are produced, what is the average fixed manufacturing cost per unit pr decimal places.) d? (Round Per unit answer to 2 f. If 5,000 units are produced, what is the total amount of fixed manufacturing cost incurred? ced, what is the total amount of manufacturing overhead cost incurred? What is this total amount expressed on a per unit basis? (Round Per unit answer to 2 decimal places. h. If the selling price is $24.30 per unit, what is the contribution margin per unit sold? (Round Perunit answer to 2 , If 3,000 units are produced, what is the total amount of direct manufacturing cost incurred? J. If 3,000 units are produced, what is the total amount of indirect manufacturing cost incurred? k. What incremental manufacturing cost will the company incur if it increases production from 4,000 to 4,001 un answer to 2 decimal places.) its? (Round Per unit
e. Average fixed manufacturing cost per unit f. Total fixed manufacturing overhead cost g. Total manufacturing overhead cost Manufacturing overhead per unit h. Contribution margin per unit i. Total direct manufacturing cost i. Total indirect manufacturing costN k. Incremental manufacturing cost
Answer #1
Saxbury Corporation
Ans e If 5000 units were produced average fixed cost per unit produced
Fixed Manufacturing overhead per unit=(A) $            4.00
Number of units budgeted=(B) 4000
Total Manufacturing cost=(C)=(A)*(B) $ 16,000.00
Number of units produced=(D) 5000
Average fixed cost per unit produced=(C )/(D) $            3.20
Ans f If 5000 units were produced total fixed manufacturing cost incurred
Fixed Manufacturing overhead per unit=(A) $            4.00
Number of units budgeted=(B) 4000
Total Manufacturing cost=(C)=(A)*(B) $ 16,000.00
Ans g If 5000 units were produced total manufacturing Overhead cost
Fixed Manufacturing overhead per unit=(A) $            4.00
Number of units budgeted=(B) 4000
Total fixed Manufacturing cost=(C)=(A)*(B) $ 16,000.00
Variable manufacturing overhead cost per unit=(D) $            1.60
Number of units produced=(E ) 5000
Total variable manufacturing cost=(F)=(D)*(E) $   8,000.00
Total manufacturing overhead cost=(G)=(C )+(F) $ 24,000.00
Total amount expressed on per unit basis
Total manufacturing overhead cost=(H) $ 24,000.00
Number of units produced=(I ) 5000
Manufacturing overhead per unit=(H)/(I) $            4.80
Ans h If selling price is $24.30 contribution margin per unit
Selling Price=(A) $         24.30
Variable Cost
Direct Material $            6.80
Direct Labor $            3.50
Variable manufacturing Overhead $            1.60
Sales commission $            0.50
Variable administrative Expenses $            0.50
Total Variable Expenses=(B) $         12.90
Contribution Margin per unit=(A)-(B) $         11.40
Ans i If 3000 units were produced total amount of direct manufacturing cost incurred
Number of units produced=(A) 3000
Direct Material cost per unit $            6.80
Direct Labor cost per unit $            3.50
Total direct manufacturing cost per unit=(B) $         10.30
Total direct manufacturing cost =(A)*(B) $ 30,900.00
Ans j If 3000 units were produced total amount of indirect manufacturing cost incurred
Number of units produced=(A) 3000
Variable Manufacturing overhead cost per unit=(B) $            1.60
Total variable cost=(C )=(A)*(B) $   4,800.00
Fixed manufacturing Cost=($4*4000)=(D) $ 16,000.00
Total Indirect manufacturing cost=(C) +(D) $ 20,800.00
Ans K Incremental cost incur if it increase production from 4000 to 4001 units
Direct Material cost per unit $            6.80
Direct Labor cost per unit $            3.50
Variable manufacturing cost per unit $            1.60
Incremental cost incur if it increase production from 4000 to 4001 units $         11.90